OpenStudy (anonymous):

Find the solutions of the given equation in the indicated interval. (Give your answers to two decimal places. Enter your answers as a comma-separated list.) 42 (cos(x))^2 + cos(x) -1 = 0 [0,pi]

OpenStudy (anonymous):

\[42(\cos (x))^2+\cos (x)-1=0\]

OpenStudy (anonymous):

That is a cleaner version of it.

myininaya (myininaya):

\[42u^2+u-1=0\] first do you know how to solve this?

myininaya (myininaya):

42(-1)=-6(7) 1=-6+7 so \[42u^2+7u-6u-1=0\] factor by grouping \[7u(6u+1)-1(6u+1)=0\] \[(6u+1)(7u-1)=0\] \[u=\frac{-1}{6}, u=\frac{1}{7}\] so we actually have \[cosx=\frac{-1}{6}, cosx=\frac{1}{7}\]

myininaya (myininaya):

\[x=\cos^{-1}(\frac{-1}{6}), x=\cos^{-1}(\frac{1}{7})\] \[x=1.7, x=1.4\]

myininaya (myininaya):

oops two decimal spots \[x=1.74, x=1.43\]